Diffusion Morning Report 09/04/07

Ted Berg tberg at stanford.edu
Tue Sep 4 06:48:36 PDT 2007


Good Morning All,
    Looking OK this Morning. . Tylan 2 is still down for H2 flow issues. 
Tylan 1 had bad uniformity.POCL3 and BBr3 still under observation. Poly 
dep rate is higher after baratron move but seems repeatable. LTO running 
ok but OT led is on. Thermco poly 1 pump is out for repair. THermco poly 
2 testing. Temp control at 300 not good on SiGe. No other problems at 
this time.

  	  	  	  	  	 
  	Date: 	9/4/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	yellow 	low dep 	  	  	 
Tylan 2 	down  	H2 flow issues 	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	yellow 	smoking 	  	  	 
Tylan 7-TEOS2 	yellow 	high pressure 	2.86/30 	  	 
Tylan 8-SiGe 	up 	  	21.64/30 	  	 
Tylan 9-Poly 	yellow 	higher dep rate 	21.98/30 	  	 
Tylan 10-Nitride 	up 	  	.3/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	ot lights 	7.29/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	down  	pump out for rebuild 	  	  	 
Thermco Poly2 	future 	needs testing 	
	
	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	A on line 540 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070904/a24b4dec/attachment.html>


More information about the diffusion mailing list